d) y=3.4**+2

d) y = 3.4*+2

Need help with the steps for solving derivative.
Answer #1

4 72+2 x²+2 (d) y = 3.4 In both sides; Taking my = m(3.4**+2) - in 3 in 4 [ Log in (a.b) = in a tlmb] = n3 + (x²+2) in a [ n(